OpenStudy (anonymous):

Can someone tell me whether this can be solved by factoring, completing the square, or using the quadratic formula. y=-x^2+10x-19

OpenStudy (kainui):

I don't think you'll be able to solve it by factoring, but you can always complete the square or use the quadratic formula. The quadratic formula is actually the "equation" version of completing the square.

hero (hero):

It can be solved by factoring.

hero (hero):

So it can be solved using all three methods.

hero (hero):

@AshleyBananas, Why did you give me the medal ?

OpenStudy (anonymous):

because you answered my question

OpenStudy (anonymous):

@Hero

hero (hero):

Did you solve it by factoring yet?

hero (hero):

@AshleyBananas, are you factoring it?

OpenStudy (anonymous):

@Hero you can't solve by factoring?

hero (hero):

Do you want me to post the steps for you?

hero (hero):

@AshleyBananas

OpenStudy (anonymous):

THANK YOU SO MUCH THAT WOULD BE GREAT! I don't understand this.

OpenStudy (anonymous):

@Hero

hero (hero):

I'll post the solution hang on a bit

hero (hero):

\[ \\-x^2 + 10x -19 \\-(x^2 - 10x + 19) \\-(x^2 +(5 + \sqrt{6})x +(5 - \sqrt{6})x - 19) \\ -(x(x + 5 + \sqrt{6}) + (5 - \sqrt{6})(x + 5 +\sqrt{6}) \\-(x + 5 - \sqrt{6})(x + 5 + \sqrt{6}) \]

OpenStudy (anonymous):

@Hero thank you so much I really appreciate it

hero (hero):

You're welcome

OpenStudy (kainui):

That's exactly how I factored in middle school.

hero (hero):

By the way, we figure it out like this: \[10 = (5 + \sqrt{6} - \sqrt{6} + 5)\] \[19 = (5 + \sqrt{6})(5 - \sqrt{6})\]

hero (hero):

Doesn't matter. Every quadratic can be factored. The question, in my opinion, is a silly one.

OpenStudy (kainui):

Obviously every one can be factored, but I think you should take into consideration the context of the course. It would seem to be implied that this means "factored into whole numbers".

OpenStudy (anonymous):

My teacher taught us that not every equation can be factored @Hero and @Kainui she does not care if they are whole numbers or not.

OpenStudy (kainui):

...

hero (hero):

I think you mean "assumptions" @Kainui. @AshleyBananas, your teacher is wrong. All you have to do is show her your work from what I posted. She may even be baffled by it.

OpenStudy (kainui):

lol ok whatever.